Regal Enterprises is considering the purchase of a new embroidering machine. It is expected to generate additional sales of RM400,000 per year. The machine will cost RM295,000, plus RM3,000 to install it. The embroiderer will save RM12,000 in labor expense each year. The embroiderer will require annual operating expenses of RM136,000. Regal is in the 34% income tax bracket. The machine will be depreciated on a straight-line basis over five years. The company plans to sell the machine at the end of year three for RM100,000.

A. Calculate the initial outlay for Regal.

B. Calculate the annual depreciation for the new embroidering machine.

C. What is the annual operating cash flow that the machine will generate?

D. Calculate terminal cash flow for Regal.
Answer #1

a) Initial Outlay = Cost of Machine+Installation Cost = 295000+3000 = RM 298000

b) Annual Depreciation = [Cost of Machine+Installment]/5 = 298000/5 = RM 59600

c) Annual Operating Cash Flow = [(Additional Sales+Savings in Labour Cost-Depreciation-Operating Expenses)*(1-Tax Rate)]+Depreciation = [(400000+12000-59600-136000)*(1-0.34)]+59600 = RM 202424

d) Terminal Cash Flow = After Tax Sale Value

Book Value after 3 years = 298000-(3*59600) = RM 119200. Therefore, Sale is at Loss.

Therefore, Terminal Cash Flow = Sale Proceeds+Tax Benefit on Loss = 100000+[(119200-100000)*34%] = 100000+6528 = RM 106528